# Integrating Multi-Preconditioned Conjugate Gradient with Additive Multigrid Strategy

### v1

@@ 3.2 Comparing components of $\boldsymbol{\alpha}$
In Example 2, we also see the similar behavior of the values of $\alpha_{\ell}$. Notably, some levels consistently exhibit negative values of $\alpha_{\ell}$. Despite this, the finest level consistently shows the largest values of $\alpha_{\ell}$ in both the blocking and conducting cases, and these values are significantly higher compared to other levels. This observation indicates that some levels in the hierarchy might not contribute positively, as suggested by the values of $\boldsymbol{\alpha}$.
@@ 4 Conclusion
In conclusion, we present a novel approach for employing the additive multigrid method in practice by exploiting the MPCG method. The numerical studies lead to the proposition that, especially in situations where an optimal hierarchy of multilevel subspaces is unavailable, the additive-MG-MPCG method can be highly beneficial. Its robust convergence properties and additive nature indicate a strong potential for outperforming the multiplicative MG-PCG method, notably in parallel computing environments. In addition to the parallel implementation of the method, we aim to extend this methodology to a matrix-free framework 9_kothari_multigrid_2022 , as both the geometric MG method and MPCG method can be implemented in a matrix-free manner. Furthermore, this work can be extended seamlessly within the unfitted finite element framework for interface problems in the domain decomposition framework 10_kothari_multigrid_2021*a .

### v2

@@ 3.2 Comparing components of $\boldsymbol{\alpha}$
In Example 2, we observe similar behavior in the values of $\alpha_{\ell}$. Notably, some levels consistently exhibit negative values of $\alpha_{\ell}$. Despite this, the finest level consistently shows the largest values of $\alpha_{\ell}$ in both the blocking and conducting cases, and these values are significantly higher compared to other levels. This observation indicates that some levels in the hierarchy might not contribute positively, as suggested by the values of $\boldsymbol{\alpha}$.
@@ 4 Conclusion
In conclusion, we present a novel approach for integrating the additive multigrid method with the MPCG method. The numerical studies lead to the proposition that, especially in situations where an optimal hierarchy of multilevel subspaces is unavailable, the additive MG-MPCG method can be highly beneficial. Its robust convergence properties and additive nature indicate a strong potential for outperforming the multiplicative MG-PCG method, notably in parallel computing environments. Future work will focus on leveraging parallel implementations, and combining the proposed methodology with a matrix-free strategy 9_kothari_multigrid_2022 . Furthermore, this work can be seamlessly extended within the unfitted finite element framework for interface problems in the domain decomposition framework 10_kothari_multigrid_2021*a .